OpenStudy (anonymous):

Simplify: fifteen raised to the negative fourth power divided by fifteen raised to the fifth power one divided by fifteen raised to the ninth power fifteen raised to the ninth power fifteen raised to the first power one-fifteenth

OpenStudy (mathstudent55):

To divide powers that have the same base, simply write the base and raise it to the difference of the exponents. Example: 28^9/28^5 = 28^(9-5) = 28^4

OpenStudy (mathstudent55):



OpenStudy (anonymous):

okay would it be the same for this? Simplify: negative nine raised to the tenth power times negative nine raised to the negative third power

OpenStudy (anonymous):

would it be (-9)^13

OpenStudy (anonymous):

is that right?

OpenStudy (mathstudent55):

When you multiply powers with the same base, you _add_ the powers, so 5^6 x 5^8 = 5^14

OpenStudy (mathstudent55):

The original problem was a division so you subtracted the powers. The new problem is a multiplication, so you add the powers.

OpenStudy (anonymous):

so it would be (-9)^13

OpenStudy (anonymous):

okay well im just going to go with that answer

OpenStudy (mathstudent55):

(-9)^10 x (-9)^(-3) = (-9)^(10 + -3) = (-9)^7